# RIS-Enabled Spoofing Against Adversary Sensing: CRB-Maximizing Design and Decoying Analysis

### v2

@@ I Introduction
In the context of the emerging paradigm of Integrated Sensing And Communications (ISAC) [9], and, particularly, in the respective concerns for sensing security and privacy [10], RISs have been lately deployed for sensing protection against adversary radars. In [11], RIS-aided stealth was studied with the goal to minimize the echo power at hostile Receivers (RXs), presenting, under idealized LoS assumptions, closed-form designs for a single monostatic radar as well as low-complexity solvers based on the mean square error criterion for multi-radar scenarios. The authors in [12] utilized a RIS to suppress returns along the true direction from a malicious monostatic radar while redirecting energy toward clutter for the radar to infer a false Angle-of-Arrival (AoA). A hybrid RIS designed was deployed in [13] to shield a region from an adversary while preserving communications and legitimate sensing. In particular, an alternating optimization strategy to jointly configure the transmit and RIS-induced beamforming was presented. Very recently, in [14], an RIS was optimized to occultate user positions from wiretappers, while maintaining both sensing and desired communication performance between the user and a legitimate Base Station (BS). However, the state of the art either relies on clutter to mislead an adversary from sensing or suppresses echo power, while there is a lack of a Cramér-Rao Bound (CRB) analysis quantifying the fundamental deception limits of RIS-based spoofing.
@@ II System and Channel Models
We consider a narrowband wireless system setup comprising an adversary monostatic $N$-antenna radar and a UE within the radar’s vicinity covered entirely by an $M$-element RIS, whose presence is unknown to the former. The radar follows a typical sensing procedure according to which, it performs UE localization via processing the received reflections of its transmitted pilot symbols. On the other hand, the RIS is optimized to obfuscate its (equivalently, UE’s) location, by synthesizing a deceptive scattering response at a prescribed decoy position, thereby, inducing false target parameter estimation, thus, masking the UE. In this initial version of the RIS spoofing framework in this paper, we consider a $2$-Dimensional (2D) geometry with half-wavelength-spaced Uniform Linear Arrays (ULAs) for both the adversary BS and the RIS. The extension to a planar RIS is left for future work.
@@ II-A Received Reflection Signal Model
$=\sum_{m=0}^{M-1}\mathbf{e}^{\jmath\omega_{n}}\,\left[\bm{\alpha}_{M}(\psi_{\text{out}})\right]_{m}^{*}\,\left[\bm{\alpha}_{M}(\psi_{\text{in}})\right]_{m},$ (2)
where $\psi_{\text{in}}$ and $\psi_{\text{out}}$ are the AoA and Angle of Departure (AoD) at and from the RIS, respectively. For an $L$-element ULA with half-wavelength antenna spacing, the steering vector of the latter kernel at an azimuth angle $\theta$ is given by:
@@ II-B Channel Model
Let $\psi_{\text{true}}$ and $\theta_{\text{true}}$ represent respectively the AoA of the BS signal at the RIS (i.e., the UE) and the AoA the reflected signal at the BS. Assuming that the BS is placed at the origin, while the RIS at the point $\mathbf{p}_{\rm RIS}\triangleq[p_{x},p_{y}]$, the round trip cascaded channel in (1) is modeled as follows:
@@ III-A CRB Analysis and Discussion
It holds from (1)’s inspection that, for any hypothesized AoA $\theta$, the received signal at the outputs of the BS’s antenna elements modeled via $\mathbf{Y}$ has the mean value $E[\mathbf{Y}]=\sqrt{P}\mathbf{g}(\theta;\bm{\omega})\mathbf{s}$ and the variance $E[(\mathbf{Y}-E[\mathbf{Y}])(\mathbf{Y}-E[\mathbf{Y}])^{\rm H}]=\sigma^{2}\mathbf{I}_{N}$, where $\mathbf{g}(\theta;\bm{\omega})\triangleq a_{\text{RIS}}\bar{\beta}(\theta;\bm{\omega})\bm{\alpha}_{N}(\theta)v(\theta)$ is the composite scalar gain comprising of this RIS kernel $\bar{\beta}(\theta;\bm{\omega})\triangleq\beta(\theta+\pi,\theta;\bm{\omega})$ and $v(\theta)\triangleq\bm{\alpha}_{N}^{\rm H}(\theta)\mathbf{f}_{\mathrm{TX}}$, are the array responses. According to [15], the scalar FI with respect to $\theta$ can be computed as follows:
@@
It is noted that, under standard array approximations, the RIS kernel changes only weakly with angle-much more slowly than the BS array response-so in a small neighborhood around the actual AoA $\theta$, the RIS kernel is typically flat, i.e., $\frac{\partial\beta(\theta;\bm{\omega})}{\partial\theta}\approx 0$ [16]. Thus, the scalar FI expression in (6) can be approximated as:
@@
Consequently, according to (8) and (9), if the RIS imposes $\bar{\beta}(\theta_{\text{true}};\bm{\omega})\approx 0$, then $J(\theta_{\text{true}};\bm{\omega})\approx 0$, implying that the CRB for $\theta_{\text{true}}$ (i.e., the error variance of the radar BS estimator) diverges and the PEB becomes unbounded at the true AoA. Conversely, concentrating the RIS kernel power at a decoy AoA $\theta_{\text{fake}}$, yields small CRB around that wrong direction.
@@ III-B Problem Formulation and Solution
Maximizing this ratio drives the leakage at $\bm{\Theta}_{\text{null}}$ to zero when feasible; otherwise, it trades off decoy angle gain against nulling efficiency over $\Theta_{\text{null}}$. Following [19, 20, 21, 22], we recast $\mathcal{P}_{1}$ under the assumptions: i) non-aliased, distinct concealment samples (i.e., $\operatorname{rank}(\mathbf{V})=K$); ii) $N\!\geq\!2K$; and iii) the decoy AoA is not included in $\Theta_{\text{null}}$ (i.e., $\mathbf{w}\notin\text{span}(\mathbf{V})$). Under these conditions a feasible solution exists and, as $\varepsilon\!\to\!0^{+}$ and $N\to\infty$, $\mathcal{P}_{1}$ converges to the same optimum value with [^3] If the hard-null constraint $\mathbf{V}^{\rm H}\bm{\omega}=\bm{0}_{M\times 1}$ is feasible, the denominator of $\mathcal{P}_{1}$ collapses to $\varepsilon$ with $\mathcal{S}$, hence, maximizing $\mathcal{P}_{1}$ is equivalent to maximizing $|\mathbf{w}^{H}\bm{\omega}|^{2}$, i.e., $\mathcal{P}_{3}$ up to a global phase configuration. On the other hand, when exact nulling is infeasible, $\mathcal{P}_{1}$ yields a smooth decoy-leakage trade-off and closely approximates the hard-null solution [23].:
$\mathcal{P}_{3}:\,\max_{\bm{\omega}}\quad|\mathbf{w}^{\rm H}\bm{\omega}|\,\,\text{s.t.}\,\mathbf{V}^{\rm H}\bm{\omega}=\bm{0}_{M\times 1},\,|[\bm{\omega}]_{m}|=1\,\forall m.$
@@ IV Analysis of the RIS-Based Decoy Capability
According to the assumptions in Section III-B, the feasible solution set $\mathcal{F}$ is compact and $\mathcal{P}_{3}$’s objective is continuous. Hence, based on the extreme value theorem [19] and [20], a maximizer $\bm{\omega}\in\mathcal{F}$ is feasible. In addition, when $N\to\infty$ with fixed $\Delta$, $J(\theta_{\text{true}};\bm{\omega})\to 0$ holds, yielding $\text{CRB}(\theta_{\text{true}};\bm{\omega})\to\infty$. In other words, any locally unbiased AoA estimator at the radar BS becomes uninformative for the true UE direction when the RIS enforces a sufficiently deep null at it. In practice, however, $|\bar{\beta}(\theta_{\text{true}};\bm{\omega})|$ becomes small but nonzero, so the CRB remains finite yet very large.
In this section, we study the conditions for $\theta_{\text{fake}}$ to effectively conceal UE’s position. Even if the hard-null constraint is satisfied when $N\!\to\!\infty$, in the finite $N$-value regime, a criterion to quantify the level of deception synthesized by the RIS is missing. To this end, we henceforth: i) introduce a novel point-wise deception criterion and ii) develop the corresponding set-wise version over $\bm{\Theta}_{\text{null}}$, as well as iii) introduce a decoy-placement score for principled selection of $\theta_{\text{fake}}$.
@@ IV-B Decoy Angle Selection Score
Expression (17) depends on $\bm{\omega}$ through $L_{\text{true}}(\bm{\omega})$. Moreover, under the feasibility conditions of $\mathcal{P}_{3}$ (i.e., $\text{rank}(\mathbf{V})=K$, $N\geq 2K$, and $\mathbf{w}\notin\operatorname{span}(\mathbf{V})$), hard nulls are achievable and $L_{\text{true}}(\bm{\omega})$ can be made arbitrarily small, hence, there is no finite upper bound on $\rho$ that holds uniformly over all $\bm{\omega}\in\mathcal{F}$. As the nulls deepen, the CRB at the true angle $\theta_{\text{true}}$ diverges. A configuration-independent upper bound is possible by imposing a design leakage cap, i.e., if $L_{\text{true}}(\bm{\omega})\ \leq\ \overline{L}$ holds for all feasible $\bm{\omega}$, then:
$\rho\ \leq\ \rho_{\text{UB}}(\theta_{\text{fake}};\bar{L})\triangleq\frac{N^{2}\,\Phi^{2}(\theta_{\text{fake}})}{\kappa_{\min}\,\overline{L}^{\,2}}.$ (18)

### v3

@@ I Introduction
In the context of the emerging paradigm of Integrated Sensing And Communications (ISAC), and, particularly, in the respective concerns for sensing security and privacy [9], RISs have been lately deployed for sensing protection against adversary radars. In [10], RIS-aided stealth was studied with the goal to minimize the echo power at hostile Receivers (RXs), presenting, under idealized LoS assumptions, closed-form designs for a single monostatic radar as well as low-complexity solvers based on the mean square error criterion for multi-radar scenarios. The authors in [11] utilized a RIS to suppress returns along the true direction from a malicious monostatic radar while redirecting energy toward clutter for the radar to infer a false Angle-of-Arrival (AoA). A hybrid RIS design was deployed in [12] to shield a region from an adversary while preserving communications and legitimate sensing. In particular, an alternating optimization strategy to jointly configure the transmit and RIS-induced beamforming was presented. Very recently, in [13], an RIS was optimized to occultate user positions from wiretappers, while maintaining both sensing and desired communication performance between the user and a legitimate Base Station (BS). However, the state of the art either relies on clutter to mislead an adversary from sensing or suppresses echo power, while there is a lack of a Cramér-Rao Bound (CRB) analysis quantifying the fundamental deception limits of RIS-based spoofing.
@@ II System and Channel Models
We consider a narrowband wireless system setup comprising an adversary monostatic $N$-antenna radar and a UE within the radar’s vicinity covered entirely by an $M$-element RIS, whose presence is unknown to the former. The radar follows a typical sensing procedure according to which, it performs UE localization via processing the received reflections of its transmitted pilot symbols. On the other hand, the RIS is optimized to obfuscate its (equivalently, UE’s) location, by synthesizing a deceptive scattering response at a prescribed decoy position, thereby, inducing false target parameter estimation, and thus, masking the UE. In this initial version of the RIS spoofing framework in this paper, we consider a $2$-Dimensional (2D) geometry with half-wavelength-spaced Uniform Linear Arrays (ULAs) for both the adversary BS and the RIS. The extension to a planar RIS is left for future work.
@@ II-A Received Reflection Signal Model
$=\sum_{m=0}^{M-1}\mathbf{e}^{\jmath\omega_{m}}\,\left[\bm{\alpha}_{M}(\psi_{\text{out}})\right]_{m}^{*}\,\left[\bm{\alpha}_{M}(\psi_{\text{in}})\right]_{m},$ (2)
where $\psi_{\text{in}}$ and $\psi_{\text{out}}$ are the AoA and Angle of Departure (AoD) at and from the RIS, respectively. For an $L$-element ULA with half-wavelength antenna spacing [14], the steering vector of the latter kernel at an azimuth angle $\theta$ is given by:
@@ II-B Channel Model
Let $\psi_{\text{true}}$ and $\theta_{\text{true}}$ represent respectively the AoA of the BS signal at the RIS (i.e., the UE) and the AoA of the reflected signal at the BS. Assuming that the BS is placed at the origin, while the RIS is at the point $\mathbf{p}_{\rm RIS}\triangleq[p_{x},p_{y}]$, the round trip cascaded channel in (1) is modeled as follows:
@@ III-A CRB Analysis and Discussion
It holds from (1)’s inspection that, for any hypothesized AoA $\theta$, the received signal at the outputs of the BS’s antenna elements modeled via $\mathbf{Y}$ has the mean value $E[\mathbf{Y}]=\sqrt{P}\mathbf{g}(\theta;\bm{\omega})\mathbf{s}$ and the variance $E[(\mathbf{Y}-E[\mathbf{Y}])(\mathbf{Y}-E[\mathbf{Y}])^{\rm H}]=\sigma^{2}\mathbf{I}_{N}$, where $\mathbf{g}(\theta;\bm{\omega})\triangleq a_{\text{RIS}}\bar{\beta}(\theta;\bm{\omega})\bm{\alpha}_{N}(\theta)v(\theta)$ is the composite scalar gain comprising of this RIS kernel $\bar{\beta}(\theta;\bm{\omega})\triangleq\beta(\theta+\pi,\theta;\bm{\omega})$ and $v(\theta)\triangleq\bm{\alpha}_{N}^{\rm H}(\theta)\mathbf{f}_{\mathrm{TX}}$, are the array responses. According to [15], the scalar FI with respect to the unknown AoA parameter $\theta$ can be computed as follows:
@@
Note that, under standard array approximations, the RIS kernel changes only weakly with angle; much more slowly than the BS array response. Thus, in a small neighborhood around the actual AoA $\theta$, the RIS kernel is typically flat, i.e., $\frac{\partial\beta(\theta;\bm{\omega})}{\partial\theta}\approx 0$ [16], indicating that the scalar FI expression in (6) can be approximated as follows:
@@
Consequently, according to (8) and (9), if the RIS imposes $\bar{\beta}(\theta_{\text{true}};\bm{\omega})\approx 0$, then $J(\theta_{\text{true}};\bm{\omega})\approx 0$, implying that the CRB for $\theta_{\text{true}}$ (i.e., the error variance of the radar BS estimator) diverges and the PEB becomes unbounded at the true AoA. Conversely, concentrating the RIS kernel power at a decoy AoA $\theta_{\text{fake}}$ yields small CRB around that wrong direction.
@@ III-B Problem Formulation and Solution
Maximizing this ratio drives the leakage at $\bm{\Theta}_{\text{null}}$ to zero when feasible; otherwise, it trades off decoy angle gain against nulling efficiency over $\Theta_{\text{null}}$. Following [19, 20, 21, 22], we recast $\mathcal{P}_{1}$ under the assumptions: i) non-aliased, distinct concealment samples (i.e., $\operatorname{rank}(\mathbf{V})=K$); ii) $M\!\geq\!2K$; and iii) the decoy AoA is not included in $\Theta_{\text{null}}$ (i.e., $\mathbf{w}\notin\text{span}(\mathbf{V})$). Under these conditions a feasible solution exists and, as $\varepsilon\!\to\!0^{+}$ and $M\to\infty$, $\mathcal{P}_{1}$ converges to the same optimum value with [^3] If the hard-null constraint $\mathbf{V}^{\rm H}\bm{\omega}=\bm{0}_{K\times 1}$ is feasible, the denominator of $\mathcal{P}_{1}$ collapses to $\varepsilon$ with $\mathcal{S}$, hence, maximizing $\mathcal{P}_{1}$ is equivalent to maximizing $|\mathbf{w}^{H}\bm{\omega}|^{2}$, i.e., $\mathcal{P}_{3}$ up to a global phase configuration. On the other hand, when exact nulling is infeasible, $\mathcal{P}_{1}$ yields a smooth decoy-leakage trade-off and closely approximates the hard-null solution [23].:
$\mathcal{P}_{3}:\,\max_{\bm{\omega}}\quad|\mathbf{w}^{\rm H}\bm{\omega}|\,\,\text{s.t.}\,\mathbf{V}^{\rm H}\bm{\omega}=\bm{0}_{K\times 1},\,|[\bm{\omega}]_{m}|=1\,\forall m.$
@@ IV Analysis of the RIS-Based Decoy Capability
According to the assumptions in Section III-B, the feasible solution set $\mathcal{F}$ is compact and $\mathcal{P}_{3}$’s objective is continuous. Hence, based on the extreme value theorem [19] and [20], a maximizer $\bm{\omega}\in\mathcal{F}$ is feasible. In addition, when $M\to\infty$ with fixed $\Delta$, $J(\theta_{\text{true}};\bm{\omega})\to 0$ holds, yielding $\text{CRB}(\theta_{\text{true}};\bm{\omega})\to\infty$. In other words, any locally unbiased AoA estimator at the radar BS becomes uninformative for the true UE direction when the RIS enforces a sufficiently deep null at it. In practice, however, $|\bar{\beta}(\theta_{\text{true}};\bm{\omega})|$ becomes small but nonzero, so the CRB remains finite yet very large.
In this section, we study the conditions for $\theta_{\text{fake}}$ to effectively conceal UE’s position. Even if the hard-null constraint is satisfied when $M\!\to\!\infty$, in the finite $M$-value regime, a criterion to quantify the level of deception synthesized by the RIS is missing. To this end, we henceforth: i) introduce a novel point-wise deception criterion and ii) develop the corresponding set-wise version over $\bm{\Theta}_{\text{null}}$, as well as iii) introduce a decoy-placement score for principled selection of $\theta_{\text{fake}}$.
@@ IV-B Decoy Angle Selection Score
Expression (17) depends on $\bm{\omega}$ through $L_{\text{true}}(\bm{\omega})$. Moreover, under the feasibility conditions of $\mathcal{P}_{3}$ (i.e., $\text{rank}(\mathbf{V})=K$, $M\geq 2K$, and $\mathbf{w}\notin\operatorname{span}(\mathbf{V})$), hard nulls are achievable and $L_{\text{true}}(\bm{\omega})$ can be made arbitrarily small, hence, there is no finite upper bound on $\rho$ that holds uniformly over all $\bm{\omega}\in\mathcal{F}$. As the nulls deepen, the CRB at the true angle $\theta_{\text{true}}$ diverges. A configuration-independent upper bound is possible by imposing a design leakage cap, i.e., if $L_{\text{true}}(\bm{\omega})\ \leq\ \overline{L}$ holds for all feasible $\bm{\omega}$, then:
$\rho\ \leq\ \rho_{\text{UB}}(\theta_{\text{fake}};\bar{L})\triangleq\frac{M^{2}\,\Phi^{2}(\theta_{\text{fake}})}{\kappa_{\min}\,\overline{L}^{\,2}}.$ (18)